OpenStudy (anonymous):

HELP , medal find the length of a missing leg of a right triangle given a leg of length 15 and a hypotenuse of length 17. Leave answer in simplest form. a. √ 225 b. 64 c. 8 d. 5 √ 5

OpenStudy (anonymous):

as 17^2=15^2 + x^2 289=225 + x^2 289-225=x^2 64 = x^2 \[\sqrt{64} = x\] 8 = x

OpenStudy (anonymous):

Oh wow thanks a lot, this really helps :)